OpenStudy (kingethan8500):

The legs of a right triangle are 3 units and 2 units. What is the length of the hypotenuse? Round your answer to the nearest hundredth.

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

do you know the pythagora's theorem ?

OpenStudy (kingethan8500):

no

OpenStudy (kingethan8500):

@kg1975

jhonyy9 (jhonyy9):

but than you not can calculi the length of hypotenuse

OpenStudy (kingethan8500):

do you know the answer?

jhonyy9 (jhonyy9):

@mathstudent55 do you have any idea different that the pythagora"s theorem ?

jhonyy9 (jhonyy9):

@welshfella

OpenStudy (kingethan8500):

@kellyspeakslouder

OpenStudy (kingethan8500):

@Mackenzie2013

OpenStudy (meowimacat):

I can help if ya want..

OpenStudy (kingethan8500):

thanks

OpenStudy (meowimacat):

So the equation for zee hypotenuse is a^2+ b^2= c^2. So plug in 2 and 3 for a nd b

OpenStudy (mathstudent55):

No. You need the Pythagorean Theorem. Otherwise, you need to use trig, and if the poster does not know the Pythagorean theorem, then there is less chance of knowing trig.

OpenStudy (mathstudent55):



OpenStudy (meowimacat):

That's what I was giving... I'm just not good at words like you are..

OpenStudy (meowimacat):

But.. yeah.. a^2+ b^2= C^2.. When you plug those numbers in, you get: 3^2+2^2= c^2

OpenStudy (meowimacat):

Are you able to do that?

OpenStudy (kingethan8500):

idk im just learning this smh

OpenStudy (kingethan8500):

is it 5.00

OpenStudy (meowimacat):

Owh, alright. Let me help you with it then I guess... So you square 3 and two (multipy it by itself). 3^2+2^2= c^2 comes out to (3*3)+(2*2)=c^2, which is essentially 9+4=c^2

OpenStudy (meowimacat):

Can you do 9+4, or should I explain that? XD nd idk what it is, i'm helping you figure it out

OpenStudy (kingethan8500):

yes

OpenStudy (meowimacat):

So.. 9+4 is equal to 13.. nd 13=c^2... but we wanna make C so it is just C, nd not c^2.. Do you know how to do that?

OpenStudy (kingethan8500):

yea i got it now thanks

OpenStudy (meowimacat):

Alrighty, you're welcome...